OpenStudy (anonymous):

Line A: 2x + 2y + 8 Line B: x + y = 4 Which statement is true about the solution to the set of equations? A. It is (8,4). B. It is (4,8). C. There is no solution D. There are infinitely many solutions

pooja195 (pooja195):

Opposites cancel -2( x + y = 4) Distribute the -2

OpenStudy (anonymous):

i dont know how to do that

pooja195 (pooja195):

-2 * x =? -2 * y =? -2 * 4 =?

OpenStudy (anonymous):

-2x, -2y, -24

pooja195 (pooja195):

-24?

OpenStudy (anonymous):

i dont know that one

pooja195 (pooja195):

2x + 2y =8 -2x -2y=-8

pooja195 (pooja195):

2-2= ?

OpenStudy (anonymous):

0

pooja195 (pooja195):

x's cancel out 2-2= ?

pooja195 (pooja195):

?

OpenStudy (anonymous):

i dont know thats hard

pooja195 (pooja195):

Are you serious?

pooja195 (pooja195):

2-2=?

OpenStudy (anonymous):

i said zero

pooja195 (pooja195):

great now 8-8= ?

OpenStudy (anonymous):

zero

pooja195 (pooja195):

good so everything cancels out therefore when we have 0=0 it is infinte solution its not always 0 it can me #=# but it has to be the same number for it to be infinte

OpenStudy (anonymous):

so its D ?

pooja195 (pooja195):

yea